OpenStudy (anonymous):

If circumference of circle R is 12 times pi is and circumference of circle S is 32 times pi , which is the ratio of a radius of S to a radius of R?

OpenStudy (anonymous):

what is the formula for the circumference of a circle?

OpenStudy (anonymous):

got the answer eh?

OpenStudy (anonymous):

2*pi*r.

OpenStudy (anonymous):

okay. what is the radius of circle R then?

OpenStudy (anonymous):

it doesnt say

OpenStudy (anonymous):

you know the formula for the circumference right? you know the actual circumference of the circle. equate them in other words, 2 pi r = 12 pi so r = ?

OpenStudy (anonymous):

1.91

OpenStudy (anonymous):

how did you get that? show me the steps.

OpenStudy (anonymous):

i did 2 times pi which gave me 6.28 then i divided that by 12

OpenStudy (anonymous):

it is not just 12, it is 12 times pi. so you have to divide 12 times pi by 2 times pi.

OpenStudy (anonymous):

ok i got 59.2

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

i did 12 times pi and 2 times pi and then divided them

OpenStudy (anonymous):

\[r = \frac{12 \times \pi}{2 \times \pi} = ?\]

OpenStudy (anonymous):

no i got 5.98

OpenStudy (anonymous):

okay. you should actually get 6. 2 goes in 12 6 times. and the pi on the numerator cancels out the pi on the denominator. so the radius is 6.

OpenStudy (anonymous):

similarly find out the radius for circle S

OpenStudy (anonymous):

i got it thnks

OpenStudy (anonymous):

great.